# Bourbaki degree of pairs of projective surfaces

### v2

@@ 5.2 Degree $6$ curves inside quadric surfaces
Assuming $\sigma$ is incompressible, as $c_{1}({\mathcal{T}}_{\sigma})=-3$, we can apply [16, Theorem 1.1] to ${\mathcal{T}}_{\sigma}$ without further restrictions, and obtain the following upper bounds for $m(\sigma)$ and $c_{3}({\mathcal{T}}_{\sigma})$:
@@
    If $4\leq m(\sigma)\leq 5$, then $c_{3}\leq 5-m(\sigma))$. In particular, when $m(\sigma)=5$, $\sigma$ is locally free;
@@
    If $0\leq m(\sigma)<4$, then $c_{3}\leq m(\sigma)^{2}-10m(\sigma)+25.$
From Section 3.1, $(d)$, we obtain:
@@
The following result is obtained using ([15], $0.1$):
@@ 5 $e\leq 1$
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 49.
From 4, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
@@
Assuming $\sigma$ is non-compressible, by 48, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
@@
gives linearly-independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
@@
Using 48, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 7 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
@@
In this case, we have equality of schemes $\Xi_{\sigma}=\operatorname{supp}(\mathcal{Q}_{\sigma})$. The Jacobian scheme has a structure of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$, which is outside the three lines, and therefore $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. The free resolution of ${\mathcal{T}}_{\sigma}$ is (obtained computationally):
@@
So we obtain $m(\sigma)=4,e=2$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=1$ corresponding to the point $p$, which is an irreducible component of $\Xi_{\sigma}$ of codimension three.
@@
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$. We have examples for the cases $c_{3}=2$ (see Example 56) and $c_{3}=4$ (see Example 57).
@@
First, the bound $c_{3}\leq 4$ is obtained from 45 for $m=m(\sigma)=3$. As we shown in 28, $e\geq 2$. We claim that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ in this case.
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [21, Proposition 1.1] for $c_{3}=0$, [3, Lemma 2.4] for $c_{3}=2$ and [14, Lemma 9.6] for $c_{3}=4$. ∎
@@
Here, the saturation of the annihilator ideal coincides with the saturation of the 0-th Fitting ideal, and the first Fitting ideal has codimension three. The support of $\mathcal{Q}_{\sigma}$ in codimension two consists of two lines $V(x_{3},x_{0})\cup V(x_{3},x_{0}-x_{1})$, and the first one has multiplicity two. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
so we get $\operatorname{Bour}(\sigma)=2,m(\sigma)=3$ and $c_{3}({\mathcal{T}}_{\sigma})=2$.
@@
$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{2}x_{3}-x_{3}^{2}&x_{3}^{2}-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})+2x_{3}(x_{1}-x_{0})\end{pmatrix}$.
Here, the 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the saturation of the first Fitting ideal is $R$. The codimension two locus of $\operatorname{supp}(\mathcal{Q}_{\sigma})$ consists of three simple lines. Thus, $m(\sigma)=3$. Moreover, $e=2$ and the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
@@
so that $\operatorname{Bour}(\sigma)=2$ and $c_{3}({\mathcal{T}}_{\sigma})=4$.
@@
Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)\leq 3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
@@
Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is stable if and only if $e>1$. From the result above, the only possibilities of $\operatorname{Bour}(\sigma)\leq 2$ are when $m(\sigma)\geq 3$, and moreover if $m(\sigma)=3$ we cannot have $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$. For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Section 4, $(a)$, $e\neq 1$. ∎
@@
If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [14, Section 9]. By the proof of [14, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
@@
From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [14, Lemma $9.3$]. If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7). From [3, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$. ∎

### v3

@@ 4.2 Degree 6 curves inside quadric surfaces
We start by applying Theorem 28 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}$, assuming $\sigma$ is incompressible. The quantities of the theorem become
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(9-6+3)\right\rceil=0$
$B$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+3c_{1}+8)\right\rceil-1=\left\lceil\frac{1}{3}(9-9+8)\right\rceil-1=1.$
Moreover, another important invariant is: $c_{2}(\mathcal{E})=7-m(\sigma)$. Direct application of the result shows the next pair of propositions:
@@
Proof.
Let us assume that $e\in[1,d-2]$ and $d\geq 3$. Then, from the Bourbaki formula and the fact that $\operatorname{Bour}(\sigma)\geq 0$, we obtain
$m(\sigma)\leq m_{0}+e(e-d)\leq m_{0}+e-d\leq m_{0}-2,$
which gives the claim. This covers the case $e=1$. If we now assume $e>1$ and apply Theorem 28 to $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a reflexive sheaf with $c_{1}=-1$, we obtain the constant is
$A$ $=\left\lceil\frac{1}{3}(c_{1}^{2}+2c_{1}+3)\right\rceil=\left\lceil\frac{1}{3}(1-2+3)\right\rceil=0.$
Hence, $5-m(\sigma)=c_{2}({\mathcal{T}}_{\sigma}(1))=c_{2}(\mathcal{E})\geq A=0$ and thus $m(\sigma)\leq 5$. ∎
@@
    If $6\leq m(\sigma)\leq 7$, then $c_{3}\leq 5-m(\sigma)$.
@@
    If $0\leq m(\sigma)<6$, then $c_{3}\leq m(\sigma)^{2}-14m(\sigma)+47$.
From Theorem A Theorem A, $(d)$, we obtain:
@@
Next, we apply Theorem 29 to the sheaf $\mathcal{E}={\mathcal{T}}_{\sigma}(1)$, a sheaf with $c_{1}=-1$ and $c_{2}(\mathcal{E})=5-m(\sigma)\geq 0$. Then, the theorem says that, for
$t$ $>\sqrt{3c_{2}+\frac{1}{4}}-\frac{3}{2}=\sqrt{15-3m(\sigma)+\frac{1}{4}}-\frac{3}{2}$
we have $h^{0}(\mathcal{E}(t))=h^{0}({\mathcal{T}}_{\sigma}(t+1))\neq 0$.
Direct application of this bound for high values of $m(\sigma)=4,5$ we get $t=0,t=1$, respectively. This shows the following proposition.
@@ 5 $e\leq 1$
To show this, we show that in any of these cases the Bourbaki degree $\operatorname{Bour}(\sigma)$ is odd for every possibility of $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, and then the result follows from 51.
From 3, we get $\operatorname{Bour}(\sigma)=e^{2}-3e+7-m(\sigma)$, so that when $m(\sigma)$ is even, $7-m(\sigma)$ is odd. We claim $e^{2}-3e$ is always an even number for $e\geq 0$ integer.
@@
Assuming $\sigma$ is non-compressible, by 50, we obtain that $e=1$. But from the formula for the Bourbaki degree with $e=1$, $m(\sigma)=5$ we obtain $\operatorname{Bour}(\sigma)=0$, and thus $\sigma$ must be free. On the other hand, if we assume ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, then $e=1$ and $\operatorname{Bour}(\sigma)=0$ give $m(\sigma)=5$. ∎
@@
gives linearly independent syzygies for $\nabla\sigma$, and thus ${\mathcal{T}}_{\sigma}\simeq\mathcal{O}_{\mathbb{P}^{3}}(-1)\oplus\mathcal{O}_{\mathbb{P}^{3}}(-2)$, with $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$ and $m(\sigma)=5$.
@@
Using 50, $e\in\{1,2\}$, and the two cases imply $\operatorname{Bour}(\sigma)=1$. Using Remark 6 for $p_{a}(B)=0,\deg(B)=1$, we obtain the $c_{3}$’s above, and both appear as examples. To conclude the last claim, we note that if $m(\sigma)\leq 3$, then
@@
The irreducible components of $\Xi_{\sigma}$ consist of three lines $V(x_{1},x_{3})$, $V(x_{1},x_{0}-x_{2})$ and $V(x_{3},x_{0}-x_{2})$ and a point $p=V(x_{2},x_{1}-x_{3},x_{0})$ outside the three lines, so $p\in\operatorname{Sing}({\mathcal{T}}_{\sigma})$. We may compute $m(\sigma)=4$, where the first line above occurs with length two and the other two occur with length one.
For $\mathfrak{p}_{1}=(x_{1},x_{3})$, we may consider $u=(x_{0})^{-1}$ and perform two elementary operations, namely $R_{2}^{\prime}=R_{2}-ux_{3}(x_{0}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite the matrix:
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{1}&1&-ux_{3}&-ux_{2}\\ x_{1}x_{3}(1-u(x_{0}-x_{2}))&0&-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u)&(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u)\end{pmatrix}$.
Thus, it sends the second basis vector $e_{2}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, and we may compute the cokernel from the second row of the remaining matrix:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3}(1-u(x_{0}-x_{2})),-x_{3}(x_{1}-x_{3}(x_{0}-x_{2})u),(x_{0}-x_{2})(x_{1}-x_{2}x_{3}u))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{1}x_{3},-x_{3}^{2},x_{1}-x_{2}x_{3}u)}$
$\simeq\kappa\langle 1,x_{3}\rangle,$
simplifying the invertible elements $(1-u(x_{0}-x_{2})),(x_{0}-x_{2})\notin\mathfrak{p}_{1}$, hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=2$.
For $\mathfrak{p}_{2}=(x_{1},x_{0}-x_{2})$, consider $u=(-x_{3})^{-1}\in R_{\mathfrak{p}_{2}}$ and perform two elementary operations $R_{2}^{\prime}=R_{2}+ux_{1}x_{3}R_{1}$ and $R_{1}^{\prime}=uR_{1}$ to rewrite $\nabla\sigma$ as
$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}ux_{1}&ux_{0}&1&-ux_{2}\\ x_{1}x_{3}(1+ux_{1})&x_{3}((x_{0}-x_{2})+ux_{0}x_{1})&0&x_{1}(x_{0}-x_{2})-ux_{1}x_{2}x_{3}\end{pmatrix}$.
Since it sends $e_{3}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
where
$f_{1}$ $=x_{1}x_{3}(1+ux_{1})$
$f_{2}$ $=x_{3}((x_{0}-x_{2})+ux_{0}x_{1})$
$f_{3}$ $=x_{1}((x_{0}-x_{2})-ux_{2}x_{3}).$
Since $f_{1}\in I$ and $x_{3}(1+ux_{1})\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{1}\in I$. From this and $f_{2}\in I$, we obtain:
$x_{3}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
and thus $I=(x_{0}-x_{2},x_{1})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$.
For $\mathfrak{p}_{3}=(x_{3},x_{0}-x_{2})$, take $u=(x_{1})^{-1}\in R_{\mathfrak{p}_{3}}$ and the elementary operations $R_{2}^{\prime}=R_{2}-ux_{1}x_{3}R_{1}$, $R_{1}^{\prime}=uR_{1}$ to obtain
$(\nabla\sigma)_{\mathfrak{p}_{3}}\sim\begin{pmatrix}1&ux_{0}&-ux_{3}&-ux_{2}\\ 0&x_{3}((x_{0}-x_{2})-ux_{0}x_{1})&x_{1}x_{3}(-1+ux_{3})&x_{1}(x_{0}-x_{2})+ux_{1}x_{2}x_{3}\end{pmatrix}$.
Since it sends $e_{1}$ to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, we may compute the cokernel using the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I},$
where
$f_{1}$ $=x_{3}((x_{0}-x_{2})-ux_{1}x_{0})$
$f_{2}$ $=x_{1}x_{3}(ux_{3}-1)$
$f_{3}$ $=x_{1}((x_{0}-x_{2})+ux_{2}x_{3}).$
Since $f_{1}\in I$ and $(x_{0}-x_{2})-ux_{1}x_{0}\in R_{\mathfrak{p}_{2}}^{\times}$, we obtain $x_{3}\in I$. From this and $f_{3}\in I$, we obtain:
$x_{1}(x_{0}-x_{2})\in I\Rightarrow(x_{0}-x_{2})\in I,$
and thus $I=(x_{3},x_{0}-x_{2})=\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$, so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa$, and thus $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. From the associativity formula over $\mathfrak{p}_{1},\mathfrak{p}_{2},\mathfrak{p}_{3}$, we then obtain
$m(\sigma)=2\cdot 1+1\cdot 1+1\cdot 1=4.$
From the Bourbaki degree formula, once we know that $e=2$, we conclude that $\operatorname{Bour}(\sigma)=1$ and this is a nearly free sequence. The complete minimal free resolution of ${\mathcal{T}}_{\sigma}$ can be obtained computationally
@@
which corresponds to the nearly free resolution given in 16. We have $c_{3}({\mathcal{T}}_{\sigma})=1$, corresponding to the point $p$, the unique irreducible component of $\Xi_{\sigma}$ of codimension three.
@@
Let $\sigma=(f,g)$ be a normal sequence with degrees $d_{f}=1$, $d_{g}=2$, such that $m(\sigma)=3$ and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable. Then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$, $\operatorname{Bour}(\sigma)=2$ and we may have $c_{3}=0,2,4$. We have examples for the cases $c_{3}=2$ (see Example 58) and $c_{3}=4$ (see Example 59).
@@
We point out that ${\mathcal{T}}_{\sigma}(1)$ will be a stable rank two reflexive sheaf of Chern classes $(-1,2,c_{3})$, since $c_{2}({\mathcal{T}}_{\sigma}(1))=5-m(\sigma)$. Then, the three possibilities $c_{3}\in\{0,2,4\}$ imply that $H^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, and therefore $e=2$. This follows from [undefw, Proposition 1.1] for $c_{3}=0$, [undefb, Lemma 2.4] for $c_{3}=2$ and [undefn, Lemma 9.6] for $c_{3}=4$. ∎
The unstable case of the above Proposition is numerically described in 27.
@@
Here, $(\Xi_{\sigma})_{1}$ has two irreducible components, the lines $V(x_{3},x_{0})$ and $V(x_{3},x_{0}-x_{1})$. Let the corresponding primes be denoted by $\mathfrak{p}_{1},\mathfrak{p}_{2}$.
Over the prime $\mathfrak{p}_{1}=(x_{0},x_{3})$, let us take $u\doteq(x_{0}-x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ and $Q\doteq x_{0}x_{1}+3x_{3}^{2}$. Consider the elementary operations $R_{2}=R_{2}-u(x_{0}x_{1}+3x_{3}^{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}ux_{3}&-ux_{3}&0&1\\ 2x_{0}x_{2}+x_{1}x_{3}-ux_{3}&x_{0}x_{3}+uQx_{3}&x_{0}^{2}&0\end{pmatrix}$,
so the matrix sends $e_{4}\in R_{\mathfrak{p}_{1}}^{4}$ to $(1,0)\in R_{\mathfrak{p}_{1}}^{2}$, so we may compute the cokernel from the second row and remaining columns:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{f_{1},f_{2},f_{3}}\doteq\frac{R_{\mathfrak{p}_{1}}}{I},$
where
$f_{1}$ $=2x_{0}x_{2}+x_{1}x_{3}-u(x_{0}x_{1}+3x_{3}^{2})x_{3}$
$f_{2}$ $=x_{3}(1+u(x_{0}x_{1}+3x_{3}^{2}))$
$f_{3}$ $=x_{0}^{2}.$
Therefore, since $f_{2}\in I$ and $u(x_{0}x_{1}+3x_{3}^{2})\in\mathfrak{p}_{1}$, we conclude $x_{3}\in I$. Rewriting $uf_{1}$ and removing terms with $x_{3}$, we obtain
$uf_{1}=2x_{0}^{3}x_{1}x_{2},$
and since $f_{3}=x_{0}^{2}\in I$, we obtain $I=(x_{0}^{2},x_{3})$, hence
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})}\simeq\kappa\langle 1,x_{0}\rangle$
so that $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=2$.
Over $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, take $u\doteq(x_{0}^{2})^{-1}$ and consider the elementary operation and $R_{2}^{\prime}=uR_{2}$, to obtain the matrix
$(\nabla\sigma)_{\mathfrak{p}_{2}}\sim\begin{pmatrix}x_{3}&-x_{3}&0&x_{0}-x_{1}\\ u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&x_{0}x_{1}+3x_{3}^{2}\end{pmatrix}$,
which sends $e_{3}\in R_{\mathfrak{p}_{2}}^{4}$ to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel using the remaining columns of the first row, namely:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}=\frac{R_{\mathfrak{p}_{2}}}{\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}}\simeq\kappa,$
and therefore $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
Thus, from the additivity formula, we obtain $m(\sigma)=3$. Moreover, $e=2$, so that $\operatorname{Bour}(\sigma)=2$.
The sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution of the form:
@@
so we get $c_{3}({\mathcal{T}}_{\sigma})=2$ and the Bourbaki scheme is a union of two skew lines.
@@
$\nabla\sigma=\begin{pmatrix}2x_{0}&2x_{1}&2x_{2}&2x_{3}\\ x_{3}(x_{2}-x_{3})&x_{3}(x_{3}-x_{2})&x_{3}(x_{0}-x_{1})&(x_{0}-x_{1})(x_{2}-2x_{3})\end{pmatrix}$.
Here, the irreducible components of $(\Xi_{\sigma})_{1}$ are three lines, given by the associated primes:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{0}-x_{1},x_{3})$
$\mathfrak{p}_{3}$ $=(x_{2}-x_{3},x_{0}-x_{1}).$
Over $\mathfrak{p}_{1}$, we may choose $u=(2x_{1})^{-1}\in R_{\mathfrak{p}_{1}}$ to perform the elementary operations $R_{2}^{\prime}=R_{2}-ux_{3}(x_{3}-x_{2})R_{1}$ and $R_{1}^{\prime}=uR_{1}$, to rewrite:
$(\nabla\sigma)_{\mathfrak{p}_{1}}\sim\begin{pmatrix}2ux_{0}&1&2ux_{2}&2ux_{3}\\ x_{3}(x_{2}-x_{3})(1+ux_{0})&0&x_{3}((x_{0}-x_{1})-2ux_{2}(x_{3}-x_{2}))&(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2})\end{pmatrix}$
Thus, it sends $e_{2}$ to $(1,0)$ and compute the cokernel by considering the remaining columns of the second row, namely
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{I}$
with $I=(f_{1},f_{2},f_{3})$ and
$f_{1}$ $=x_{3}(x_{2}-x_{3})$
$f_{2}$ $=x_{3}((x_{0}-x_{1})-2ux_{3}(x_{3}-x_{2}))$
$f_{3}$ $=(x_{0}-x_{1})(x_{2}-2x_{3})-2ux_{3}^{2}(x_{3}-x_{2}).$
Since $f_{1}\in I$, $x_{3}(x_{2}-x_{3})\in I$, and we may rewrite $f_{2}$ as
$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\in I\Rightarrow x_{3}(x_{0}-x_{1})\in I,$
and therefore $x_{3}\in I$. Now, turning to $f_{3}\in I$ and removing terms with $x_{3}$, we obtain:
$f_{3}\equiv x_{2}(x_{0}-x_{1})\Rightarrow x_{2}\in I,$
since $x_{0}-x_{1}$ is invertible. Thus, $I\simeq\mathfrak{p}_{1}R_{\mathfrak{p}_{1}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}=1$.
Over $\mathfrak{p}_{2}=((x_{0}-x_{1}),x_{3})$, we may choose the same elementary operations and thus
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\frac{R_{\mathfrak{p}_{2}}}{I}$
with $I=(f_{1},f_{2},f_{3})$, the polynomials as before. Since $f_{1}\in I$, we conclude:
$f_{3}=(x_{0}-x_{1})(x_{2}-2x_{3})+2ux_{3}^{2}(x_{2}-x_{3})\Rightarrow(x_{0}-x_{1})(x_{2}-2x_{3})\in I.$
Since $(x_{2}-2x_{3})$ is invertible in $R_{\mathfrak{p}_{2}}$, we obtain $(x_{0}-x_{1})\in I$. From this fact, since $x_{2}-x_{3}$ is invertible and $f_{1}\in R_{\mathfrak{p}_{2}}$, we obtain $x_{3}\in I$, and thus $I\simeq\mathfrak{p}_{2}R_{\mathfrak{p}_{2}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$.
Over the prime $\mathfrak{p}_{3}$, we may also do the same elementary operations and consider the same generators for the ideal $I$. From $f_{1}\in I$ we get $(x_{2}-x_{3})\in I$, since $x_{3}$ is now invertible. Moreover, from $f_{2}\in I$ we obtain:
$f_{2}=x_{3}(x_{0}-x_{1})+2ux_{2}x_{3}(x_{2}-x_{3})\Rightarrow x_{3}(x_{0}-x_{1})\in I\Rightarrow(x_{0}-x_{1})\in I,$
so that $I\simeq\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$ and $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}=1$. Thus, from the associativity formula $m(\sigma)=3$. Moreover, $e=2$ and thus $\operatorname{Bour}(\sigma)=2$.
Here, the sheaf ${\mathcal{T}}_{\sigma}$ admits a free resolution below
@@
so that $c_{3}({\mathcal{T}}_{\sigma})=4$, and the associated Bourbaki scheme is a plane smooth quadric.
@@
Let $\sigma$ be a normal sequence with $d_{f}=1,d_{g}=2$. If $m(\sigma)<3$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
@@
Since $\mu({\mathcal{T}}_{\sigma})=-3/2$, then ${\mathcal{T}}_{\sigma}$ is $\mu$-stable if and only if $e>1$. For $m(\sigma)<3$, $\operatorname{Bour}(\sigma)>2$ and by Theorem B Theorem B, $(a)$, $e\neq 1$. ∎
@@
If $c_{3}=9$, then $E={\mathcal{T}}_{\sigma}(1)$ has Chern classes (1,3,9)(-1,3,9), thus it is an extremal sheaf in the sense of [undefn, Section 9]. By the proof of [undefn, Lemma 9.3] we obtain $h^{0}(E(1))=2$, and therefore $h^{0}({\mathcal{T}}_{\sigma}(2))\neq 0$, thus $e=2$.
@@
From the resolution, we compute $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, a contradiction since $h^{0}({\mathcal{T}}_{\sigma}(2))=2$ from [undefn, Lemma $9.3$]. If $c_{3}=7$, then ${\mathcal{T}}_{\sigma}(1)$ is stable with Chern classes (1,3,7)(-1,3,7). From [undefb, Theorem $3.15$], we have the cohomology table of ${\mathcal{T}}_{\sigma}(1)$, and in particular $h^{0}({\mathcal{T}}_{\sigma}(2))=1$, so that $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=2$ and $p_{a}(B)=0$. ∎